# Spin precession effects in the phasing formula of eccentric compact binary inspirals till the second post-Newtonian order

### v1

@@ II.5 Spin effects in the Fourier domain phase (SUA)
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{294955\nu}{2728}-\frac{294955}{5456}\biggr)\kappa_{A}+\frac{35057425\delta}{87274176}-\frac{294955\delta\kappa_{S}}{5456}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{294955\delta\kappa_{A}}{10912}-\frac{3137401645\nu}{87274176}+\biggl(\frac{294955\nu}{5456}-\frac{294955}{10912}\biggr)\kappa_{S}+\frac{35057425}{174548352}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{294955\delta\kappa_{A}}{32736}+\frac{26037775\nu}{29091392}+\biggl(\frac{294955}{32736}-\frac{294955\nu}{16368}\biggr)\kappa_{S}-\frac{26037775}{58182784}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{294955}{16368}$
$-\frac{294955\nu}{8184}\biggr)\kappa_{A}-\frac{26037775\delta}{29091392}+\frac{294955\delta\kappa_{S}}{16368}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{294955\delta\kappa_{A}}{32736}+\frac{26037775\nu}{29091392}+\biggl(\frac{294955}{32736}$
$-\frac{294955\nu}{16368}\biggr)\kappa_{S}-\frac{26037775}{58182784}\biggr]\biggr\}+\biggl(\frac{y_{0}}{y}\biggr)^{50/3}\biggl[\biggl(\frac{77002025\nu}{2663184704}-\frac{77002025}{5326369408}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}$
$-\frac{77002025\delta}{2663184704}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{77002025}{5326369408}-\frac{77002025\nu}{2663184704}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}+\frac{77002025\delta}{2663184704}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}$
$+\biggl(\frac{77002025\nu}{2663184704}-\frac{77002025}{5326369408}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+\biggl(\frac{77002025}{5326369408}-\frac{77002025\nu}{2663184704}\biggr)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]$ (25f)
III Comparison with exact eccentricity and implications on GW data analysis
@@ III.2 Comparison with exact eccentricity TaylorT2
Figure 1: Absolute value of the difference in the number of GW cycles between a numerical evolution of the TaylorT2 phase (valid for any initial eccentricity $e_{0}$) and the analytical TaylorT2 phase in Eq. (21). The 1-cycle line roughly represents where the phase error between the exact eccentricity solution and the analytical low eccentricity solution becomes significant and our approximation breaks down. The above plot has been made for a system with component masses $m_{1}=m_{2}=1.4\,M_{\odot}$. The magnitude of the individual spins are chosen to be $\bm{\chi}_{1}||=0.7$, $\bm{\chi}_{2}||=0.8$, with their orientations being on the x-y plane. Note that the total angular momentum vector is directed along the z-axis.
@@ IV Conclusions and Discussions
In this work, we have developed—for the first time—closed-form analytical expressions that consistently capture the joint influence of orbital eccentricity and generic spin precession on the gravitational-wave (GW) phasing of compact binaries up to second post-Newtonian order. By employing the precession-averaging method of [77], we effectively removed the explicit time dependence in the spin–orbit and spin–spin interactions, thereby enabling tractable analytic solutions to the eccentricity evolution. Treating eccentricity as a small parameter, we obtained phasing formulae accurate up to the eighth power of the initial eccentricity $e_{0}$, and provided both time-domain (TaylorT2) and frequency-domain (TaylorF2/SUA) representations. Furthermore, we proposed a simple resummation procedure that extends the domain of validity of our TaylorT2 phase to eccentricities as high as $e_{0}\sim 0.8$, thereby considerably enhancing the practical applicability of our results for data analysis.
@@ Acknowledgements.
We are particularly grateful to Gonzalo Morras for clearing a significant number of queries during the course of this study, and Divyajyoti for assistance with the review process. We thank Patricia Schmidt for useful comments and suggestions on our manuscript. This document has LIGO preprint number LIGO-P2500628.

### v2

@@ II.5 Spin effects in the Fourier domain phase (SUA)
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{294955\delta\kappa_{A}}{32736}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{864463355\nu}{43637088}}+\biggl(\frac{294955}{32736}-\frac{294955\nu}{16368}\biggr)\kappa_{S}-\frac{26037775}{58182784}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{294955}{16368}-\frac{294955\nu}{8184}\biggr)\kappa_{A}-\frac{26037775\delta}{29091392}+\frac{294955\delta\kappa_{S}}{16368}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{294955\delta\kappa_{A}}{32736}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{294955\nu}{16368}}+\biggl(\frac{294955}{32736}-\frac{294955\nu}{16368}\biggr)\kappa_{S}-\frac{26037775}{58182784}\biggr]\biggr\}+\biggl(\frac{y_{0}}{y}\biggr)^{50/3}\biggl[\biggl({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{77002025\nu}{1331592352}}$
$-\frac{77002025}{5326369408}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}-\frac{77002025\delta}{2663184704}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{77002025}{5326369408}$-
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{77002025\nu}{1331592352}}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}+\frac{77002025\delta}{2663184704}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{77002025}{5326369408}}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{77002025}{5326369408}}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]$ (26b)
Note that these coefficients up to $\mathcal{O}(e_{0}^{8})$ are given in [104].
III Comparison with exact eccentricity and implications for GW data analysis
@@ III.2 Comparison with exact eccentricity TaylorT2
Figure 1: Absolute value of the difference in the number of GW cycles between a numerical evolution of the TaylorT2 phase (valid for any initial eccentricity $e_{0}$) and the analytical TaylorT2 phase in Eqs. (19) and (22). The 1-cycle line roughly represents where the phase error between the exact eccentricity solution and the analytical low eccentricity solution becomes significant and our approximation breaks down. The above plot has been made for a system with component masses $m_{1}=m_{2}=1.4\,M_{\odot}$. The magnitude of the individual spins are chosen to be $\bm{\chi}_{1}||=0.7$, $\bm{\chi}_{2}||=0.8$, with their orientations being on the x-y plane. Note that the total angular momentum vector is directed along the z-axis.
III.3 Mismatch computations
To estimate the effects of the spin-precessing eccentric corrections to the waveform, we first included the newly computed analytic frequency domain phasing to a private branch of LALSuite. The phasing correction includes the eccentric spin-precessing terms until 2 PN, as well as eccentric aligned-spins terms until 3 PN. We however, include only the Newtonian amplitude to our waveform and restrict the initial eccentricity accuracy up to $\mathcal{O}(e_{0}^{2})$. Then we used the standard definition of match as noise weighted inner product of two waveforms, one of which contains our newly computed phasing terms. We define the noise weighted inner product as
$\left(h_{1}|h_{2}\right)\equiv 4\Re\left[\int_{f_{start}}^{f_{end}}\frac{\tilde{h}_{1}(f)\tilde{h}_{2}^{*}(f)}{S_{n}(f)}\right]$ (31)
where $\tilde{h}_{1}(f)$ and $\tilde{h}_{1}(f)$ are the Fourier transforms of $h_{1}$ and $h_{2}$ respectively, $f_{start}$ and $f_{end}$ are the start and end cutoff frequencies which we take to be $20$ Hz and a fraction of the corresponding Kerr ISCO frequency ($0.8\times f_{Kerr,ISCO}$) respectively, $S_{n}(f)$ is the power spectral density (PSD), and * denotes complex conjugation. We take $S_{n}(f)$ to be the advanced LIGO, zero detuned, and high laser power configuration. We define the match as the following [108]
$\mathcal{M}(h_{1},h_{2})\equiv\,\dfrac{(h_{1}|h_{2})}{\sqrt{(h_{1}|h_{1})(h_{2}|h_{2})}}\,$, (32)
such that $\mathcal{M}\mathcal{M}=1$-$\mathcal{M}$ is the mismatch between the two waveforms $h_{1}$ and $h_{2}$.
We first perform some benchmark studies of our waveform PrecTaylorF2Ecc with TaylorF2Ecc, which contains circular spins and is accurate up to $\mathcal{O}(e_{0}^{2})$. As a first test, we compute the mismatch of PrecTaylorF2Ecc with TaylorF2Ecc while varying the purely perpendicular initial spin configurations, as is defined in the co-precessing frame with respect to the unit orbital angular momentum vector, and for an equal mass system for two values of eccentricity in Fig. 2.
Figure 2: In the above figures, we compute the mismatch between TaylorF2Ecc and PrecTaylorF2Ecc for two different values of initial eccentricity $e_{0}$. In the left figure we used a contour plot to quantify the mismatch for various initial x component of spins $\chi_{1,x}$ and $\chi_{2,x}$ for zero initial eccentricity. While on the right figure, we plot the same for $e_{0}=0.3$. One notices that while the trend is almost the same for both the figures, however, the center of the figures reveal a slight increase in the mismatch value when the initial eccentricity is increased from zero to 0.3.
We then perform a mismatch computation between PrecTaylorF2Ecc and pyEFPE, first by varying the chirp mass ($\mathcal{M}_{chirp}$) and $e_{0}$ for fixed values of an arbitrary spin configuration in the left plot of Fig. 3, and then by varying the effective spin precession parameter $\chi_{p}$, as can be found in Ref. [109], and $e_{0}$, for an equal mass and z-component of spins in the right plot of Fig. 3.
Figure 3: In the left figure, we set the initial spin configuration as $\chi_{1,x}=0.6$, $\chi_{2,x}=0.7$, $\chi_{1,z}=\chi_{2,z}=0.2$, varied the chirp mass and initial eccentricity, and plotted the mismatch for various values of the same between PrecTaylorF2Ecc and pyEFPE. As can be seen, for lower chirp masses which are of long duration, PrecTaylorF2Ecc accumulates more error given that the initial eccentricity accuracy order is $\mathcal{O}(e_{0}^{2})$, and that it uses a Newtonian amplitude. However, the mismatch improves, and hence the usability is better for higher chirp masses. We also plot the $5\%$ and $10\%$ mismatch regions and notice that PrecTaylorF2Ecc performs best around $20\,M_{\odot}$. In the right figure, we fix the masses to be $20\,M_{\odot}$, $\chi_{1,z}=\chi_{2,z}=0.2$, and varied the effective spin precession parameter $\chi_{p}$ and the initial eccentricity. We find that PrecTaylorF2Ecc performs better in the small eccentricity and small precession limit, whereas it performs poorly for high eccentricities and high precession. We also plot the $1\%$ and $2\%$ mismatch regions to illustrate the domain of validity of PrecTaylorF2Ecc.
@@ IV Conclusions and Discussions
In this work, we have developed, for the first time, closed-form analytical expressions that consistently capture the joint influence of orbital eccentricity and generic spin precession on the gravitational-wave (GW) phasing of compact binaries up to second post-Newtonian order. By employing the precession-averaging method of [87], we effectively removed the explicit time dependence in the spin–orbit and spin–spin interactions, thereby enabling tractable analytic solutions to the eccentricity evolution. Treating eccentricity as a small parameter, we obtained phasing formulae accurate up to the eighth power of the initial eccentricity $e_{0}$, and provided both time-domain (TaylorT2) and frequency-domain (TaylorF2/SUA) representations. Furthermore, we proposed a simple resummation procedure that extends the domain of validity of our TaylorT2 phase to eccentricities as high as $e_{0}\sim 0.8$, thereby considerably enhancing the practical applicability of our results for data analysis.
@@
A frequency domain Taylor waveform was also developed along with this study, which we call as PrecTaylorF2Ecc. In this waveform, developed in a private branch of LALSuite, we included our newly calculated eccentric and spin-precessing phasing terms up to 2 PN, as well as spin-aligned eccentric phasing terms up to 3 PN. Benchmark tests indicate a near perfect match with TaylorF2Ecc for zero spins, and high mismatch with the inclusion of spins. Similarly, a waveform comparison test with pyEFPE revealed the domain of validity of PrecTaylorF2Ecc in various parameter spaces. We found that our waveform is valid for low eccentricity, high chirp mass, and low $\chi_{p}$, which is expected since our TaylorF2 waveform only includes eccentricity corrections at the leading order, with no corrections beyond the Newtonian order (zero eccentricity) for the amplitude. Inclusion of higher order eccentricity terms in the TaylorF2 waveform will improve the scenario drastically in the future, which however, is beyond the scope of the present work.
@@ Acknowledgements.
We are particularly grateful to Gonzalo Morras for clearing a significant number of queries during the course of this study, Divyajyoti for assistance with the review process, and Kaushik Paul for assistance with the mismatch analysis. We thank Patricia Schmidt for useful comments and suggestions on our manuscript. This document has LIGO preprint number LIGO-P2500628.